### v1

# Tenspiler: A Verified Lifting-Based Compiler for Tensor Operations

Abstract
Tensor processing infrastructures such as deep learning frameworks and specialized hardware accelerators have revolutionized how computationally intensive code from domains such as deep learning and image processing is executed and optimized. These infrastructures provide powerful and expressive abstractions while ensuring high performance. However, to utilize them, code must be written specifically using the APIs / ISAs of such software frameworks or hardware accelerators. Importantly, given the fast pace of innovation in these domains, code written today quickly becomes legacy as new frameworks and accelerators are developed, and migrating such legacy code manually is a considerable effort.
To enable developers in leveraging such DSLs while preserving their current programming paradigm, we describe our experience in building Tenspiler, a verified lifting-based compiler that uses program synthesis to translate sequential programs written in general-purpose programming languages (e.g., C++ or Python code that does not leverage any specialized framework or accelerator) into tensor operations. Central to Tenspiler is our carefully crafted yet simple intermediate language, named Tensir, that expresses tensor operations. Tensir enables efficient lifting, verification, and code generation. Unlike classical pattern-matching-based compilers, Tenspiler uses program synthesis to translate input code into Tensir, which is then compiled to the target API / ISA. Currently, Tenspiler already supports six
DSLs, spanning a broad spectrum of software and hardware environments. Furthermore, we show that new backends can be easily supported by Tenspiler by adding simple pattern-matching rules for Tensir. Using 10 real-world code benchmark suites, our experimental evaluation shows that by translating code to be executed on 6 different software frameworks and hardware devices, Tenspiler offers on average 105$\times$ kernel and 9.65$\times$ end-to-end execution time improvement over the fully-optimized sequential implementation of the same benchmarks.
keywords:
Program Synthesis, Code Transpilation, Tensor DSLs, Verification
1 Introduction
We have witnessed an explosion of new computational infrastructures for tensor computation in recent years: from software frameworks such as TensorFlow to specialized hardware accelerators like tensor processing units. Such infrastructures arise due to new application domains such as image processing and training deep learning (DL) models, and they often expose their functionality via various domain-specific languages (DSLs) that range from specialized instruction sets such as vectorized instructions to high-level programming interfaces such as Apple’s MLX [28] or Tensorflow’s XLA [35].
To leverage the optimization offered by such infrastructures, applications must be written against the provided programming interfaces: developers must first master each DSL’s programming model to write new code, and existing applications must be rewritten. This problem is recurring as new DSLs keep appearing targeting different application domains. Manually rewriting existing applications is tedious and increases the likelihood of introducing bugs. The classical way of addressing such issues is to build transpilers [15, 5, 6, 20, 26, 27] that translate code from paradigms developers are familiar with (e.g., C++ code using the STL library) to the one provided by the target DSL (e.g., NumPy API). Nonetheless, building such a transpiler is resource-intensive, error-prone, and each one is specialized to a specific target DSL. For instance, existing compilers such as Dexter [6], STNG [20], and C2Taco [26] target specific DSLs like Halide and TACO, and are not easily extensible to support new operators or backends. While recently developed DL models such as GPT have shown promise in code translation, they do not provide any guarantees on the correctness of output. Moreover, GPT fails to generate even syntactically correct code for DSLs it has not seen in training data, limiting its applicability to new or less popular DSLs.
In this paper, we describe our experience in building a tensor compiler that addresses these challenges. We introduce Tenspiler—a compiler designed to automate the transpilation of code to multiple tensor processing frameworks and hardware accelerators. Tenspiler uses verified lifting [12] (VL), a technique using inductive program synthesis to infer provably equivalent program summaries expressed using a user-defined intermediate representation (IR), and generate executable code from the synthesized summary to the target DSL. In contrast to conventional compilers that rely on pattern-matching to compile code, VL uses a search-based technique for the translation process. The two key steps of VL are:
  • Search Phase: This stage lifts the input code to an equivalent program written using a user-provided IR, where the IR is used to model the functionality of each operator in the target DSL. Lifting is formulated as a syntax-guided synthesis [8] problem.
  • Verification: Once lifted, a theorem prover is used to validate if the synthesized summary is functionally equivalent to the input. If so, executable code is produced by calling the user-provided code generator from the summary, otherwise, another summary is generated by the search phase.
The key to making lifting efficient lies in the design of the IR (i.e., how the target DSL is modeled). In prior work [5, 6, 7], each function or instruction exposed by the target DSL is modeled explicitly. While doing so makes the search efficient, such explicit modeling makes the compiler hard to extend to other DSLs. With Tenspiler, we introduce, for the first time, a single unified IR, Tensir, that is designed for tensor operations and can easily generate code to multiple tensor processing software frameworks and hardware accelerators. Surprisingly, Tensir is a small language based on tensor algebra that includes commonly used vector and matrix operations. While other unifying IR exists (e.g., MLIR [24]), they are targeted for classical pattern-matching compilers. As we will discuss in Sec. 5 and Section 4.2, Tensir instead is designed for synthesis-based compilers and thus aims to enable both efficient search and verification.
In summary, this experience paper makes the following contributions:
  1. 1.
    We describe the design of Tensir for transpiling code to tensor processing DSLs. Tensir is simple yet expressive enough to model the functionalities provided by different software frameworks and hardware accelerators, and enables efficient code transpilation using verified lifting, as detailed in Sec. 3.
  2. 2.
    Based on Tensir, we devise various optimization techniques to make synthesis and verification tractable, and scale to real-world programs in Sec. 5.
  3. 3.
    We implemented Tenspiler, a verified lifting-driven transpiler built using Tensir as the modeling language. We demonstrate the effectiveness of Tenspiler by using it to lift real-world code from 10 different suites to 6 different open-source and commercially-available tensor processing software frameworks and hardware accelerators. We demonstrate the ease of constructing such transpilers by building one for MLX, a new tensor processing framework that was released only four months ago, using less than 200 lines of code in Sec. 6.
2 Overview
1inline uint8_t screen_8x8 (uint8_t a, uint8_t b) { return a + b - (a * b) / 255; }
2vector<vector<int>> screen_blend(vector<vector<int>> b, vector<vector<int>> a) {
3 vector<vector<int>> out; int m = b.size(); int n = b[0].size();
4 for (int row = 0; row < m; row++) {
5 vector<int> r_v;
6 for (int col = 0; col < n; col++)
7 r_v.push_back(screen_8x8(b(col, row), a(col, row)));
8 out.push_back(r_v);}
9 return out;}
(a) Original Blend function in C++.
1def t_t(x, y, operation):
2 if len(x) < 1 or len(x) != len(y): return []
3 else: return [operation(x[0], y[0])] + t_t(x[1:], y[1:], operation)
4
5def t_s(x, a, operation):
6 if len(x) < 1: return []
7 else: return [operation(x[0]), a] + t_s(x[1:], a, operation)
(b) Operators in Tensir. We represent tensor_scalar as t_s and tensor_tensor as t_t.
1def inner_loop(row, col, b, a, r_v, out):
2 return col >= 0 and col <= len(b[0]) and row >= 0 and row < len(b) and
3 r_v == t_t(t_t(b[row][:col], a[row][:col], +),
4 t_s(t_t(b[row][:col], a[row][:col], *), 255, /), -) and
5 out == t_t(t_t(b[:row], a[:row], +), t_s(t_t(b[:row], a[:row], *), 255, /), -)
6
7def outer_loop(row, col, b, a, row_vec, out):
8 return row >= 0 and row < len(b) and
9 out == t_t(t_t(b[:row], a[:row], +), t_s(t_t(b[:row], a[:row], *), 255, /), -)
(c) Synthesized loop invariants.
1def screen_blend(b, a): return b + a - b * a // 255 #NumPy/TensorFlow/PyTorch/MLX
2uchar256 Screen8x8(uchar256 a, uchar256 b) { #TPC-C implementation for Gaudi
3 uchar256 c = v_u8_mul_b(a, b) * v_reciprocal_fast_f32(255);
4 uchar256 d = v_u8_add_b(a, v_u8_sub_b(b, c));
5 return d; }
(d) Generated executable code for different tensor processing DSL.
Figure 1: End-to-End example of using Tenspiler to transpile code
Tenspiler takes in C/C++ or Python code as input [^1] Tenspiler currently supports a subset of the C/C++ and Python language (in particular it does not support code that uses pointers or objects, which we have not encountered such use in our benchmarks). It also expects any external libraries used in the input to be functionally modeled, which is how Tenspiler currently supports code that uses the STL::vector library. and transpiles it to a functionally equivalent program that leverages different software frameworks and hardware accelerators (details described in  Section 4.3) for tensor computation. As mentioned, Tenspiler uses verified lifting to first translate the source program into Tensir. Unlike traditional pattern-matching compilers, Tenspiler formulates code translation as a search for a program expressed in Tensir that is provably semantic-equivalent to the input. Doing so avoids the need to devise pattern-matching rules and prove their correctness. To make the search scalable, instead of directly searching within the DSL exposed by each target, we designed a high-level IR called Tensir that abstracts away the low-level implementation details of each DSL operator and captures only their semantics, unifying various DSLs into a common set of tensor operators.
Tenspiler uses a program synthesizer (currently Rosette [34], a synthesizer for finite domain theories) to lift the input code to Tensir during the search phase. The synthesized output is then verified using a theorem prover (currently, an SMT solver, CVC5 [10]) for the unbounded domain. “Unbounded domain” means the verification is performed for all possible program states, not just a bounded set of states (e.g., all states where integers are represented using 8 bits) that Rosette considers during the synthesis phase.
Once verified, Tenspiler then translates the Tensir program to the concrete syntax of the target DSL using simple pattern-matching rules.
We illustrate Tenspiler using the example in Figure 1(a) as our S (source), where S implements blending, a common image processing operation. It lightens the base color by iterating over each pixel, implemented as a nested loop over all the rows and cols in the image. Our goal is to transpile this code to the target DSLs supported by Tenspiler as shown in Figure 1(d).
[…]
and is defined recursively on each element. Meanwhile, t_s
performs element-wise scalar operations on tensor x using the scalar value a and is similarly defined. Importantly, both operators are purely functional models of the tensor operations that lack implementation details that a specific target might leverage (e.g., tiling, vectorization, etc). The idea is that if S can be expressed using only these operators via lifting, then the lifted program can be easily translated to the targeted backends.
In Tenspiler, lifting is formulated as a Syntax-Guided Synthesis (SyGuS) [8] problem, where the goal is to synthesize a semantically equivalent program summary (PS), represented as a sequence of operators from our Tensir, with the input code as the specification. A search space (specified using grammar) describes the set of potential candidate programs for the given specification. An input program S is semantically equivalent to the synthesized expression S’ if for all possible program inputs i, S(i) = S’(i).
Tenspiler uses symbolic search to solve the synthesis problem. Symbolic search is typically implemented through enumerative or deductive search, and using constraint-solving approaches which often rely on domain-specific heuristics to scale.
As a SyGuS problem, symbolic search is implemented as enumerating different expressions over a user-provided grammar, where the grammar encodes all possible combinations of operators in the target DSL up to a specified depth. However, as the depth increases, the number of choices grows exponentially, making the search intractable
As we will discuss in Sec. 5, Tensir is designed to make synthesis scalable. For S, the synthesis phase returns the following solution:
def lifted_program(b, a): return t_t(t_t(b, a, +), t_s(t_t(b, a, *), 255, /), -)
As Tenspiler’s synthesizer currently can only reason about finite domains, all synthesized solutions are checked for full functional equivalence using an automated theorem prover. Since S has loops, checking equivalence with the generated program on all inputs requires loop invariants. Such invariants are synthesized during the synthesis phase by constructing a grammar similar to the PS grammar. For instance, for S, the synthesis phase yields two loop invariants (one for each loop) alongside a PS. As shown in Figure 1(c), these loop invariants are not arbitrary; within the loop invariants, the output variable, out, is expressed as a combination of operators from the Tensir that help prove the synthesized PS. We will leverage this to improve synthesis efficiency, to be explained in Sec. 5
With the synthesized solution expressed in Tensir, the final step is to translate it into the concrete syntax of the target DSL(s). In Tenspiler, this is done via simple pattern-matching rules. In Figure 1(d), we present the translated code for different supported DSLs. As we will discuss in Section 4.3, Tensir is designed such that generating executable code is straightforward. In fact, the code generators for the different tensor processing infrastructures supported by Tenspiler are highly similar to each other, as we will discuss in Sec. 6.
3 The Tensir Intermediate Representation
We now discuss our intermediate representation, Tensir, which plays a pivotal role in Tenspiler. While prior lifting-based compilers all use search to compile programs, their IRs are specialized for their use cases. Those IRs consist of operators from the target languages, describing their high-level semantics while avoiding low-level implementation details. For example, Casper [5] was built to translate sequential Java to MapReduce programs. The MapReduce framework consists of several versions of map that differ by their input types. However, Casper only defines one operator in its IR that models map’s functionality, and decides on the implementation to use during code generation. While doing so makes synthesis tractable, it also makes the compiler inflexible as adding another target (e.g., a hardware accelerator that supports map over tensors) will require modeling its functionality, which may be incompatible with the existing map from Casper’s IR.
To address this challenge, we designed a novel IR, Tensir, by studying the DSLs provided by various software frameworks and hardware accelerators for tensor computation. Tensir is rooted in tensor algebra and is designed for flexibility, allowing translation to both software (deep learning frameworks, vector processing libraries) and hardware environments (machine learning accelerators) as to be discussed in Section 4.3. This flexibility enables developers to select which target to execute the translated code based on availability and specific performance requirements. Given the dynamic nature of tensor processing infrastructures, Tensir can be modified easily in terms of both adding support for new tensor operators and new target backends. This is illustrated in Sec. 6, where it only took 200 lines of code for Tenspiler to support Apple’s recently introduced MLX framework [28].
Comparison with MLIR. MLIR is a compiler infrastructure that enables the representation and transformation of code at various levels of abstraction. The core idea behind MLIR is to provide a unified IR that can capture the semantics of the program at different levels of detail (dialects), from high-level abstractions down to low-level, target-specific instructions. Developers can use MLIR by progressively lowering the code through different dialects until it reaches a level suitable for the target hardware. While MLIR and its dialects offer a powerful infrastructure for progressively targeting multiple hardware backends, we found that the existing dialects do not fully support all the operators required for our use case. Independently, both the linalg and tensor dialects do not support all the operators Tensir supports. For example, the $select$ operator, which is crucial for image processing kernels that apply operations conditioned on pixel values, is not supported by any MLIR dialects. Additionally, unifying these dialects can be challenging for developers. Instead of unifying, recent work such as mlirSynth [14] has explored using program synthesis to translate between different MLIR dialects. In contrast, Tensir is designed to be flexible and easily extensible. Developers can add new operators to Tensir by simply describing their high-level semantics, and new backend support can be incorporated by defining simple pattern-matching rules. This approach allows developers to extend Tensir without going into the intricacies of MLIR. Moreover, Tensir can practically be compiled into different MLIR dialects, providing developers the flexibility to leverage the MLIR infrastructure if desired.
3.1 Language Definition
The operators and grammar of Tensir are shown in Fig. 2. Tensir operates on tensors [^2] In the grammar, the $Tensor\;Literal$ refers to 1D or 2D tensors as we did not encounter higher dimensional tensors in our benchmarks. and includes various operations. The core strength of Tensir lies in tensorOp, which forms the backbone of tensor operations, including a diverse range of manipulations on tensors, such as element-wise operations, tensor vector multiplication, and reductions. These are grouped into different categories:
  • tensor_scalar operations describe element-wise operations involving tensors and scalars, such as scalar multiplication of each element in a matrix.
  • tensor_tensor operations perform element-wise operations between two tensors, such as element-wise multiplication of two tensors.
  • Tensor reshaping such as transpose.
  • tensor_vec_prod operation denotes tensor-vector products, enabling operations like matrix-vector multiplication. [^3] While we can also define a tensor-tensor product operator, we did not encounter such benchmarks in our evaluation and hence omitted it from Tensir’s grammar.
  • Tensor reductions such as reduce_max and reduce_sum, which focus on aggregating tensor values, with the former determining the maximum element and the latter computing the sum across specified dimensions.
$p\in Op$ ::=:= $s_{comp}\;\mid t_{comp}\;\mid\;c_{control}$
$t_{comp}\in tensorOp$ ::=:= $tensor\_scalar(t,l,o)\;\mid\;tensor\_tensor(t,t,o)\;\mid\;$
$transpose(t)\;\mid\;tensor\_vec\_prod(t,t)\;\mid\;a$
$s_{comp}\in scalarOp$ ::=:= $reduce\_max(t)\;\mid\;reduce\_sum(t)$
[…]
$c_{control}\in controlflowOp$ ::=:= $ite(cond,i,i)$
$cond\in boolExpr$ ::=:= $i\;rop\;i$
$i\in inp$ ::=:= $l\;\mid\;t$
[…]
$a\in accessOp$ ::=:= $take(t,l)\;\mid\;tail(t,l)\;\mid\;slice(t,d,l_{1},l_{2})$
$l$ ::=:= $Integer\;Literal\;\mid\;size(t)\;\mid\;s_{comp}$
$t$ ::=:= $Tensor\;Literal\;\mid\;t_{comp}$
Figure 2: Tensir grammar
The recursive nature of Tensir’s grammar allows tensor operations to be composed, facilitating the expression of complex algorithms encountered in source code. Tensir extends its expressiveness beyond tensor operations by also including a control flow operator (controlflowOp)
It integrates control flow through the ite operator, enabling conditional logic into tensor computation. This operator is crucial for translating real-world loopy programs that contain branches.
Tensir is notable not only for its diversity of operations but also for the granularity of each operator, which significantly enhances its utility in translating code. The fine-grained nature of operations, from basic element-wise computations to advanced tensor reductions and control flow constructs, allows the grammar to capture the diverse tensor computation present in the input. Moreover, the selected set of operations aligns with the core functionalities supported by most tensor processing infrastructures. This ensures that Tensir can seamlessly integrate with various frameworks and accelerators, offering flexibility in supporting multiple target DSLs. This comprehensive yet concise grammar serves as a bridge between traditional loop-based programming paradigms and the highly parallelizable world of tensor computation, providing a clear and expressive language for describing mathematical operations on tensors.
Besides tensor operations, Tensir also supports tensor accessing and manipulation:
  • take(t, n) extracts n elements from the beginning.
  • tail(t, n) returns all the elements in tensor t after the first n elements.
  • slice(t, d, s, e) extracts a contiguous sub-tensor from t from indices s (inclusive) to e (exclusive) along dimension d.
  • size(t) returns the total number of elements in tensor t.
Such functions are used to express the loop invariants and program summaries in the synthesis phase, as we describe next.
4 Transpiling Code Using Tenspiler
Figure 3: An overview of the Tenspiler Framework.
As shown in Fig. 3, Tenspiler is designed to translate a program in high-level languages, source (S),
into another program that leverages different tensor processing infrastructures. Tenspiler currently support a vector processing library (NumPy) for CPU execution, DL frameworks (PyTorch, TensorFlow, MLX) for GPU execution, and ISAs for specialized hardware accelerators (Gaudi, Gemmini). Tenspiler is a verified-lifting-based compiler, leveraging search to find a program within the target domain. Instead of relying on traditional pattern-matching rules, Tenspiler translates source programs with a 3-phase workflow: 1. synthesis, 2. verification, and 3. backend code generation. Tenspiler uses a single IR, Tensir, to facilitate all 3 phases of its workflow. Tensir is designed to include tensor processing operators common to all target backends. As shown in the figure, the synthesis phase takes in S and generates a program summary expressed using Tensir. Then, in the verification phase, Tenspiler verifies the generated summaries to ensure their semantic equivalence with S. Finally, in the code generation phase, the Tensir program is translated to the concrete syntax of the target DSL(s).
4.1 Synthesis
The objective of this phase is to search for a program expressed using the operators in Tensir, and to ensure that the generated program is semantically equivalent to S. We formulate the search as a SyGuS problem [8] characterized by three parameters: 1. the specification describing the property the synthesized Tensir expression should satisfy, 2. the search space that describes the space of possible solutions, 3. the search algorithm which searches for the candidate programs.
For Tenspiler, the specification is to find a functionally equivalent program to S. Various methods exist to express this specification, such as using input-output examples, bounded-model checking, and verification conditions (VC) [19]. In Tenspiler, we use VCs as the specification for the synthesis phase as it provides full guarantees (i.e., for all program states up to a bound, e.g., states where all integers are encoded using 8 bits)
on the equivalence of S and the translated program. VCs are logical expressions encoding the correctness properties of S. Specifically, given a program $P$ with $vars$ representing all the variables appearing in $P$, and $pre$, $post$, $inv$ representing the pre-conditions, the post-condition, and the invariant(s), respectively, the VCs for a program with loops consist of the following clauses:
  1. 1.
    Initial Condition: $\forall\;vars.\;pre(vars)\rightarrow inv(vars)$: loop invariants must hold before the loop begins its execution.
  2. 2.
    Loop Preservation: $\forall\;vars,vars^{\prime}.\;inv(vars)\;\land\;P(vars,vars^{\prime})\rightarrow inv(vars^{\prime})$: if the invariant holds before a loop iteration, they should continue to hold after that iteration.
  3. 3.
    Post-Condition: $\forall\;vars.\;inv(vars)\rightarrow post(vars)$: invariants should hold once the loop has completed its execution.
There exists standard techniques for generating VCs from a given source program [11]. In Tenspiler, the PS and invariants in the VCs are generated as placeholders as S is analyzed,
with their bodies to be synthesized during the synthesis phase.
Next, we define the search space for synthesis. This space outlines the potential solutions for both the PS and invariant(s), describing the solutions that could potentially satisfy the VC. Expressed as a context-free grammar (CFG), the search space imposes syntactic constraints on the structure of the outputs. In Tenspiler, the goal is not to find any PS or invariants but ones that represent the output variables in S as some sequence of operators from Tensir, expressed mathematically as:
[…]
This states that all return variables in S should be expressed as a program from Tensir. With the specification in the form of VCs, $p$ expressed using Tensir, and the search space for the PS and invariants, the synthesis problem can be formally defined as:
$\exists\;inv_{0},\;inv_{1},\;\ldots,\;PS\in G.\;\forall\sigma.\;VC(S,\;inv_{0},\;inv_{1},\;...,\;PS,\sigma)$ (2)
The goal of synthesis is then to find expressions from the search space $G$ for $PS$ and $inv$s such that, for all program states $\sigma$, they satisfy the VC. The synthesizer’s effectiveness in solving Eq. 2 depends on the search space’s description, as an overly expressive grammar can degrade synthesis performance. Conversely, a restrictive grammar may cause synthesis to fail in finding suitable expressions.
Various approaches exist for implementing the search to solve Eq. 2, from symbolic approaches like enumeration, deductive search, or constraint solving, to neural-based methods. For Tenspiler, we use an off-the-shelf symbolic search engine, Rosette [34], which uses constraint solving to address the synthesis problem. In a constraint solving-based approach, the specification $\phi$ (i.e., VC) and the search space $G$ are encoded as a single formula, and an SMT solver is then utilized to find a model that satisfies the formula. As a constraint-based solver, increasing the number of constraints makes the problem more challenging. Given that $\phi$ is fixed for a particular benchmark, the design of the $G$ becomes crucial. In Sec. 5, we discuss how the design of Tensir helps keep the grammar size reasonable and scales the synthesis process.
4.2 Verification
During the synthesis phase, as the PS and loop invariant(s) are validated against the VC only for a bounded set of program states, [^4] We are unaware of any SyGuS solvers that can validate against an unbounded set of program states efficiently, including state of the art solvers such as Z3 and CVC5. it is essential to check their validity for all program states. Tenspiler uses an SMT solver to do so by negating the VC in program verification i.e., checking if $\;\neg VC(S,\;inv_{1},\;inv_{2},\;...,\;PS,\;\sigma)$
is satisfiable for some $\sigma$. The placeholders in the VC are substituted with the synthesized bodies of $PS$ and $inv$s. If the solver cannot find any such $\sigma$, then the generated $PS$ and $inv$s are correct for all possible program states, thus proving $PS$ and $inv$s hold for all program states. If a $\sigma$ is found, then Tenspiler will iterate back to the synthesis phase in search of another candidate expression.
Besides using SMT solvers for Eq. 2, Tenspiler also leverages SMT solvers’ support of algebraic data types (ADT) to allow users to define common data structures such as lists and tuples. Internally, Tenspiler models tensors using the list data structure defined using ADTs. We use ADT’s accessor and constructor functions to retrieve and create new tensors. All the tensor accessing functions like slice, take are modeled as recursive functions over the list data structure.
Currently, while image processing kernels use integers and deep learning kernels operate over floats, we verify all the benchmarks using the theory of integers and reals, due to poor solver support for reasoning about floats.
Since the verification of loop invariants is undecidable in general, we define additional axioms for the operators in Tensir to aid verification. These axioms describe the behavior of functions that cannot be automatically deduced by the solver. Identifying the axioms requires an understanding of the program’s semantics and the properties that need verification. Such axioms describe simple attributes such as distributivity, associativity, and commutativity of the tensor operators. In Fig. 4, we show an inductive axiom for the tensor_scalar operator which states that the result for a given index is determined by the product of the first element of the tensor and an integer, plus the result for the remaining sub-tensor up to that index. As shown, having tensors as first-class objects in Tensir greatly simplifies the task of defining these properties. Instead of defining these properties using low-level SMT-LIB list data structures, Tensir enables users to define them at the tensor level, abstracting away the low-level solver-related details. This high-level representation greatly simplifies the task of defining these properties and makes the axioms more readable and maintainable.
1(assert (forall ((data (Tensor Int)) (a Int) (idx Int))
2(=> (>= index 0) $\land$ (<= index len(data))
3(= t_s(data[:idx],a,*) (+ [data[0]*a] t_s(data[1:idx], a, *))))))
Figure 4: Example of an inductive axiom for the tensor_scalar operator in Tensir described using SMT-LIB. “+” corresponds to the concat operator.
4.3 Code Generation
After successful verifying the synthesized Tensir program, the final stage in Tenspiler’s workflow is to translate the Tensir program into the concrete syntax of a target DSL. Tensir makes this easy as it inherently represents tensor operations supported by all the target DSLs, and code can be generated using simple syntax-driven translation from Tensir operators to their DSL-specific counterparts.
1# Generating code for TPC-C language of Intel’s Gaudi processor
2gaudi_translations = {
3 "vec_elemwise_add": lambda args: f"v_u8_add_b({codegen(args[0],gaudi_translations)},{codegen(args[1],gaudi_translations)})",
4 ... }
5# Generating code for TensorFlow/PyTorch/MLX/NumPy DSLs
6tf_pt_mlx_np_translations = {
7 "vec_elemwise_add": lambda args: f"{codegen(args[0],tf_pt_mlx_np_translations)}+{codegen(args[1]),tf_pt_mlx_np_translations}",
8 ... }
9def codegen(expr: Expr, translations: Dict[str, Any]) -> str:
10 if isinstance(expr, CallExpr): return translations[expr.name()](expr.args)
11 ...
Figure 5: Code generation for the element-wise add operator to different targets.
Fig. 5 shows part of Tenspiler’s code generator for vec_elemwise_add operator, which is a recursive function that takes as input a Tensir expression and translates every occurrence of vec_elemwise_add into the corresponding operation in the target DSL. For TPC-C [22], which is C-based programming language supported by Intel’s Gaudi processor, it generates v_u8_add_b intrinsic. [^5] To be precise, our code generator generates v_u8_add_b intrinsic for unsigned chars, while for float32 vectors it would generate v_f32_add_b. Meanwhile, it generates the + operator for TensorFlow, PyTorch, MLX, and NumPy.
This direct and syntactic translation simplifies the integration of new tensor-based target DSL into Tenspiler, as one would only need to add simple translation rules in the code generation process. For instance, we add support for MLX by adding only 65 lines of code to an existing 130-line template, as its API closely follows that of NumPy.
As a part of this work, we have implemented support for six different target DSLs in our code generator: NumPy, TensorFlow, PyTorch, MLX (an ML framework for Apple silicon), TPC-C (C-based programming language for Intel’s Gaudi processor), and Gemmini (an open-source neural network accelerator generator). We provide further details of these DSLs, including the sizes of all the benchmark programs translated in those DSLs by Tenspiler, in Appendix A.
5 Synthesis Optimizations
$out$ ::=:= $m_{4}\;\mid\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v_{4}}\;\mid\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}l_{4}}$
$m_{4}$ ::=:= ${\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}tensor\_scalar(m_{3},\;l_{4},\;o)}\;\mid\;tensor\_tensor(m_{3},\;m_{3},\;o)\;\mid\;$
$transpose(m_{3})\;\mid\;{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}ite(cond_{4},\;m_{3},\;m_{3})}\;\mid\;m_{3}$
$v_{4}$ ::=:= $tensor\_scalar(v_{3},\;l_{4},\;o)\;\mid\;tensor\_tensor(v_{3},\;v_{3},\;o)\;\mid\;$
$tensor\_vec\_prod(m_{3},v_{3})\;\mid\;ite(cond_{4},\;v_{3},\;v_{3})\;\mid\;v_{3}$
$l_{4}$ ::=:= $reduce\_max(l_{3})\;\mid\;reduce\_sum(l_{3})\;\mid\;l_{3}$
$cond_{4}$ ::=:= $l_{3}\;rop\;l_{3}$
$l_{1}$ ::=:= $255\;\mid\;size(t_{1})$
$t_{1}$ ::=:= $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}{a\langle a_{1},\;a_{2}\;...\;a_{n}\rangle}\;\mid\;\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}{b\langle b_{1},\;b_{2}\;...\;b_{n}\rangle}\;\;\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{a\langle a_{1},\;a_{2}\;\rangle}\;\;\mid\;\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{b\langle b_{1},\;b_{2}\rangle}$
[…]
Figure 6: A depth 4 general synthesis grammar for the source in Figure 1(a)
A naive approach to constructing the grammar for search space is to enumerate all possible combinations of Tensir expressions up to a fixed depth. For Tensir as defined in Fig. 2, if we focus solely on the compute operators, a depth-4 grammar (i.e., sequence of 4 operators) results in a search space of $\sim$200k expressions, since it grows exponentially with the depth and the number of operations. In Fig. 6, we show a small part of the depth-4 grammar. We have devised several optimizations to reduce the search space and make the search tractable.
5.1 Restricting Operators
[…]
all reduction operations are excluded. In Fig. 6, the operators in $v_{4}$ and $l_{4}$ will be removed (shown in red). These correspond to operators returning 1-D vectors and integers respectively.
5.2 Restricting Program States
We further optimize the search space by restricting the set of program states in Eq. 2. Instead of satisfying the VC for all $\sigma$, we find $ps$ and $inv$ that satisfy a bounded set. Bounded synthesis is crucial because most SyGuS solvers have limited support for recursive function definitions and require SMT solvers for validation. However, SMT solvers lack inherent support for reasoning about Tensir operators that are not covered by the standard theories defined in SMT-LIB [9] and require additional axioms to be defined. We instead integrate bounded synthesis by restricting the maximum unrollings of recursive operators, thereby eliminating the need for additional axioms. Specifically, we restrict the program states by limiting the lengths of the data structures and the sizes of the data types. For instance, in Tenspiler, we constrain all 1D tensors to length 2 and the integers to 6 bits or less for the first rounds of synthesis. In Fig. 6, all the tensor literals in $t_{1}$ are changed from an unbounded length “n” (shown in orange) to length 2 (shown in blue). If the synthesized choices fail to verify, we then increase the bounds in subsequent rounds. Note that since the synthesized solutions only work for a restricted set of program states, we invoke the theorem prover for subsequent verification to check if $ps$ and $inv$ are valid for all states.
5.3 Leveraging Expression Trees
Despite the above two optimizations, the synthesis search space remains large. For example, in the context of S in Figure 1(a) for which we need to synthesize two $inv$s and one $ps$, a depth-4 grammar, after removing the reduction operations, still presents around 100k potential solutions just for $ps$. Tensir plays a significant role in the further pruning of this search space. The design of Tensir operators effectively bridges the gap between high-level tensor operations and the loop-based paradigm commonly used for computing on tensors. This property of Tensir allows us to leverage the expression tree-based filtering approach, which we describe next, to efficiently prune the synthesis search space.
Our approach starts with the static analysis of S to identify the computations performed; the static analysis pass emits an expression tree that represents the computation. For example, the pre-order traversal of the expression tree for S from Figure 1(a) is: (- (+ b a) (/ (* b a) 255)). In Fig. 6, this results in the pruning of tensor_scalar and ite(cond5, m4, m4) at the top-level (shown in teal) and similarly operators at other depths ($m_{4},m_{3}$) are filtered. The generated expression tree is then transformed into an abstract expression tree, where variables and constants are replaced with placeholders, resulting in a synthesis template.
The abstract expression tree for S is then: (- (+ var var) (/ (* var var) lit)). This abstract expression tree guides Tenspiler in identifying the sequence of Tensir operators. In this example, Tenspiler deduces the sequence of operators from the tree as: t_t(t_t(var, var, +), t_s(t_t(var, var, *), lit, /), -), where var and lit are variables and literals to be synthesized, respectively.
Our expression trees are amenable to vectorized operations, which simultaneously perform the same computation on multiple data elements. Specifically, each level of the tree corresponds to an operation with the branches indicating data flow. In the example expression shown above, we see element-wise subtraction, addition, scalar division, and element-wise multiplication orchestrated such that it aligns with the vectorized execution of the original computation.
This approach is not confined to specific operators but is adaptable to a range of operations in Tensir. It can identify constructs like if-else blocks, where $ite$ arguments are determined using the same expression tree strategy, allowing the synthesis process to determine the optimal sequence of operators within the constructs. This flexibility extends to reduction operators and other complex operations, aiding in the synthesis of efficient operational sequences.
5.4 Constraining Variables
The final optimization is to
pinpoint specific variables (vars such as a,b in Figure 1(a)) and literals (lits such as 255 in Figure 1(a)) to be used in the grammar. Specifically, we constrain the variables to the set of live variables and also constrain constants to the set of constants that have appeared in the program. This strategy simplifies the computational task, avoiding the complexity of synthesizing a complete depth-4 operator sequence. By leveraging our expression tree-based approach, the synthesizer promptly yields the correct solution within 76 secs: t_t(t_t(b, a, +), t_s(t_t(b, a, *), 255, /), -).
5.5 Overall Synthesis Algorithm
The algorithm described in Fig. 7 summarizes the synthesis phase in Tenspiler. This phase is used to search for the bodies of PS and invariants which satisfy the VC. The synthesis is an iterative process conducted over multiple rounds, assuming a filtered search space leveraging type-based and expression tree
optimizations described earlier. We start with the tensor bound size set to 2 (line 2) which corresponds to the restricting the program states optimization. In each round, we invoke Rosette’s search algorithm (line 5) to generate candidates for $ps$ and $invs$. Upon obtaining a solution, the candidate undergoes validation against the VC for all program states, as the synthesis phase only checks within a constrained set of program states. We invoke a verifier (CVC5) (line 9) to perform this check. If the verifier yields “UNSAT”, the generated candidates are correct. Conversely, if “SAT” is returned, indicating incorrect candidates (line 11), the VC is augmented with blocking constraints. These constraints state that the generated $ps$ or $inv$ in next round should differ from those in the previous rounds. This iterative process continues for a specified number of rounds (max_rounds) before incrementing the tensor bound sizes. In cases where Rosette’s search algorithm does not produce a solution initially (line 13), indicating an overly restrictive grammar, the initial grammar is expanded to include additional options for both PS and invariants, such as choices for loop bounds, indexing, and operator sequences. We keep a separate timer (not shown in Fig. 7) that maintains a maximum time bound for the entire synthesis process.
1def synthesis_algorithm(spec, holing_grammar, search_algorithm, verifier, max_rounds, timeout):
2 r, k = 0, 2 # rounds within one list bound, tensor size bound
3 #bounded synthesis optimization
4 while r < max_rounds:
5 ps_inv = search_algorithm(spec, holing_grammar) #rosette
6 if ps_inv is not None:
7 ps_r, inv_r = ps_inv
8 if verifier(specification, ps, inv) == "UNSAT": return ps, inv
9 else:
10 spec = spec and (ps != ps_r) and (inv != inv_r) #add blocking constraint
11 r += 1
12 else:
13 expand_holing_grammar(holing_grammar)
14 k += 1
15 return None # No solution found
Figure 7: Tenspiler synthesis algorithm.
6 Experiments
We evaluate Tenspiler’s effectiveness in converting code into various tensor processing infrastructures using 10 loop-based real-world benchmark sets: blend, Llama [25], blas, darknet, dsp, dspstone, makespeare, mathfu, simpl_array, and utdsp. The blend benchmarks focus on image processing kernels, the Llama benchmarks contain traditional deep learning applications, and the rest 8 are all sourced from various existing software libraries, such as the BLAS linear algebra library and the TI signal processing library, and are used recently to evaluate C to TACO translations  [26]. This combination of benchmarks allows for a comprehensive assessment of Tenspiler’s effectiveness and adaptability across diverse domains and programming paradigms.
  1. 1.
    Used in prior work [6], the blend benchmarks consist of 12 functions dedicated to point-wise image blending operations — a fundamental aspect of image processing known for diverse visual effects. These functions span 180 lines of C++ code, and 10 are characterized by doubly-nested loops, which are common in image processing algorithms.
  2. 2.
    Llama benchmarks are derived from the C++ based inference code of Llama2 [25], an open-source LLM from Meta. We labeled the portion of code to be lifted from the source code without doing any extensive syntax or code logic edits. These benchmarks include 11 functions capturing essential operations like computing activations, attention mechanisms, and layer norms. They total around 106 lines of code, with 2 functions incorporating doubly-nested loops.
  3. 3.
    blas: 2 benchmarks from the BLAS [13] linear algebra library.
  4. 4.
    darknet: 10 neural network functions sourced from the Darknet [2] deep learning framework.
  5. 5.
    dsp: 12 signal processing functions from the TI library [4].
  6. 6.
    dspstone: Kernels from the DSPStone suites [36] that target digital signal architectures.
  7. 7.
    makespeare: Programs originally from Rosin [30] that manipulate integer arrays.
  8. 8.
    mathfu: 11 mathematical functions extracted from the Mathfu library [3].
  9. 9.
    simpl_array: 5 functions for computations on integer arrays originally from the work of So and Oh [33].
  10. 10.
    utdsp: Kernels from the UTDSP suite [32] that targets digital signal architectures.
6.1 Evaluation Setup
The synthesis and verification phases for all benchmarks are executed on a MacBook Pro 2 GHz Quad-Core Intel Core i5 Processor with a timeout of 1 hour. After lifting the code to Tensir, the code generator, as explained in Section 4.3, generates executable code for each target backend. In the next section, we first describe the datasets used for evaluating the performance of lifted benchmarks. Then, we describe each target backend used for executing these benchmarks.
6.1.1 Datasets for Evaluation
To mimic real-world settings, we evaluate the translated code on actual datasets instead of generating random inputs.
For the blend image processing benchmarks, blas, darknet, dsp, dspstone, makespeare, mathfu, simpl_array, and utdsp, we source images from ImageNet [17], a large-scale image dataset. We process these images as grayscale to ensure pixel values fall within the appropriate range. For benchmarks with 1-D tensor inputs, we flatten the images before feeding them as inputs and pass them in as they are for 2-D tensor inputs. For the blending layers in the blend benchmarks, we generate random pixel values from 0 to 255. We randomly select a set of 10,000 images from the dataset for evaluation.
For the Llama benchmarks, we evaluate the synthesized code using weights from Vicuna [16], an open-source LLM with similar model size as Llama2666We did not use Llama2 model weights as they are not publicly available.. Some kernels operate over model weights, for which we directly use the weight matrices from Vicuna. For kernels operating over inputs, we simulate embeddings by creating random 32-bit float vectors within the range [0, 1). The evaluation primarily uses the 33B-parameter Vicuna model; however, for evaluating the MLX framework, the 7B-parameter version is used due to memory limitations.
6.1.2 Target Software Frameworks and Hardware Accelerators
The core objective of Tenspiler is to translate sequential programs to a spectrum of diverse target DSLs, which can then be executed on conventional CPUs, GPUs, or specialized hardware accelerators. Although finding the optimal target DSL for the given input program would be an interesting feature in Tenspiler, currently Tenspiler is designed to provide users with the flexibility of choosing their preferred environment.
For our experimental evaluation, we choose 6 different target DSLs as we mentioned in Section 4.3: NumPy, TensorFlow, PyTorch, MLX, TPC-C, and Gemmini. We believe that our comprehensive selection of DSLs is necessary to test the robustness of Tenspiler.
The Tensir design greatly simplified this process as NumPy, TensorFlow, PyTorch, and MLX have similar APIs, and each of these DSLs uses only 200 lines of code for generating executable code.
To establish a baseline for execution performance, we compile C++ code for all the benchmarks using gcc-8.3.0 with -O3 flag and then run them on an Intel Xeon 8380 CPU. Given that each DSL is tailored to enhance performance on specific hardware, we conduct evaluation across five distinct platforms: the Intel Xeon 8380 CPU, Nvidia V100 GPU, Apple M1 Pro, Intel Gaudi2 processor, and the Gemmini accelerator [^7] Due to lack of physical device, Gemmini evaluations are done on a simulator with limited computing power and no file system support. Thus, it is compared with smaller random inputs.. In all, we utilized 7 different DSL-hardware device combinations for our experiments: (1) NumPy-CPU, (2) TensorFlow-V100, (3) PyTorch-V100, (4) MLX-Apple M1 (5) TPCC-Gaudi, (6) PyTorch-Gaudi, and (7) Gemmini. This comprehensive mapping of each backend to its corresponding device enables us to accurately assess the benefits Tenspiler provided through lifting.
The exact configurations of all the hardware devices and their software environments are described in the (Appendix C) due to lack of space.
6.2 Synthesis Timings
In this section, we present the time Tenspiler takes to synthesize and verify solutions for each of our benchmarks. During the synthesis phase, we apply all the optimizations mentioned in Sec. 5 with a 1-hour timeout. Tenspiler synthesizes the correct translations for all the benchmarks under 15 mins.
(a) Synthesis timings for the blend benchmarks (image processing)
(b) Synthesis timings for the utdsp benchmarks
(c) Synthesis timings for the DSP benchmarks
(d) Synthesis timings for the simpl_array benchmarks
(e) Synthesis timings for the darknet benchmarks
(f) Synthesis timings for the blas benchmarks
(g) Synthesis timings for the mathfu benchmarks
(h) Synthesis timings for the dspstone benchmarks
(i) Synthesis timings for the Llama benchmarks (ML kernels)
(j) Synthesis timings for the makespeare benchmarks
Figure 8: Synthesis timings for all the benchmark suites. Benchmark name legend in  Appendix B.
Fig. 8 illustrates the synthesis performance for our 10 benchmark suites. Figure 8(a) illustrates the synthesis performance for the blend benchmarks. All but three benchmarks are synthesized in one synthesis (and verification) iteration, with an average synthesis time of 40.7 seconds and a median synthesis time of 2.357 seconds. Single-loop benchmarks synthesize with an average of 2.17 seconds and a median of 1.92 seconds. Double-loop benchmarks, on the other hand, have an average of 128.91 seconds and a median of 22.74 seconds for synthesis.
The three benchmarks that take more than one round of synthesis are softmax1, transformer1, and transformer2 from the Llama suite. softmax1 fails to synthesize within one round because the initial grammar is overly restrictive for its loop invariant. transformer1 and transformer2 involve complex indexing constraints for their invariants, initially leading to spurious solutions with tensor size limit of 2. transformer2 finds the correct solution after 6 tries, while transformer1 exhausts the maximum number of tries (10) with tensor size 2. We then increase the tensor size limit to 3 for transformer1 and a correct solution is generated within 3 rounds.
6.2.1 Analysis
We recognize that the tree approach may restrict the ability of Tenspiler to synthesize different solutions. However, through manual verification, we confirm that Tenspiler consistently generates optimal solutions across our benchmarks. We evaluate the solutions using expression length as the cost function. Generally, shorter expressions mean fewer function calls and thus lower execution cost. To illustrate, we use the linear_dodge example from the blend benchmarks for which the synthesized solution is as follows:
def linear_dodge(a, b): t_t(a, b, +)
After relaxing the constraint on the tree structure, we were able to synthesize a different solution as shown below:
def linear_dodge(a, b): t_t(a, t_s(b, -1, *), -)
The latter solution is longer in length, and involves 2 tensor operations – a tensor_tensor operation and a tensor_scalar operation – as opposed to 1 tensor_tensor operation synthesized using the tree approach. Therefore, it is less optimized. In addition, the tree approach also speeds up the synthesis process as shorter expressions are easier to synthesize.
(a) blend benchmarks.
(b) utdsp benchmarks.
(c) dsp benchmarks.
(d) simpl_array benchmarks.
(e) darknet benchmarks.
(f) blas benchmarks.
(g) mathfu benchmarks.
(h) dspstone benchmarks.
(i) Llama benchmarks.
(j) makespeare benchmarks.
Figure 9: Kernel speedup over baseline. Benchmark legend in  Appendix B.
6.3 Performance Timings
In this section, we evaluate the performance of the original input code—sequential C++ programs compiled with gcc -O3 —by comparing them with their translated versions executed across different target backends as detailed in Sections 6.1.2 and 6.1.1.
Kernel Performance. Kernel performance focuses on computation time excluding data transfer overhead. We see significant improvements as illustrated in Figures 9, with an average speedup of 105.1$\times$ across all benchmarks. Notably, the Gaudi2 processor demonstrates an exceptional speedup of 241$\times$, highlighting the advantages of migrating legacy code to newer hardware platforms. For other backends such TensorFlow, PyTorch, MLX and NumPy we see speedups of 46$\times$, 244$\times$, 10.5$\times$ and 26.12$\times$ respectively.
However, compatibility issues can emerge with certain backends. For example, the Gemmini accelerator does not support certain operations in our Tensir like tensor_tensor element-wise multiplication, slice, and tail. To address this, we only translate supported Tensir operations from the synthesized PS, and default to running the unsupported operations using sequential C on CPUs. Out of 69 benchmarks, 41 are translatable to Gemmini’s instruction set architecture (ISA), yet only 10 can be fully expressed using Gemmini instructions alone. Challenges are notable in benchmarks like screen_blend (Figure 9(a)), where element-wise vector multiplication must fallback to execution on a less powerful CPU. Furthermore, most Gemmini’s instructions require square matrices inputs. This means that we need to pad vector inputs to square matrices before being able to utilize Gemmini’s instructions, effectively squaring the data volume to be processed. This results in varied performance as shown in Figure 9(i) and Figure 9(e).
(a) blend benchmarks.
(b) utdsp benchmarks.
(c) dsp benchmarks.
(d) simpl_array benchmarks.
(e) darknet benchmarks.
(f) blas benchmarks.
(g) mathfu benchmarks.
(h) dspstone benchmarks.
(i) Llama benchmarks.
(j) makespeare benchmarks.
Figure 10: E2E speedup over baseline. Benchmark legend in  Appendix B.
End-to-end Performance. While frameworks and accelerators deliver substantial kernel performance enhancements, a comprehensive assessment must account for end-to-end benchmark times, encompassing initial setup and data movement between the host (CPU) and the accelerator device. Our focus here is on data transfer (TensorFlow, PyTorch, and Gaudi processor) and memory management (C++). As illustrated in  Fig. 10, we again observe an overall speedup, averaging 9.7$\times$. In particular, CPU libraries like NumPy and MLX show more improvements with the notable advantage of avoiding transferring data to specialized hardware. These benchmarks, involving the processing of 1D or 2D character vectors, benefit largely from C++’s efficient handling of contiguous data structures. Meanwhile, Gaudi2 drivers encounter performance bottlenecks due to the overheads associated with hardware initialization and frequent small data transfers. This significant upfront cost, especially pronounced in small-scale data operations, leads to a much less announced speedup. We believe such a phenomenon is uncommon in real-world use cases such as training deep learning models, due to techniques like batch processing or pipelining to minimize data transfers or to overlap computations with communications, thereby reducing or hiding transfer overhead and enhancing overall efficiency.
6.4 Ablation Study
(a) blend benchmarks.
(b) utdsp benchmarks.
(c) dsp benchmarks.
(d) simpl_array benchmarks.
(e) darknet benchmarks.
(f) blas benchmarks.
(g) mathfu benchmarks.
(h) dspstone benchmarks.
(i) llama benchmarks.
(j) makespeare benchmarks.
Figure 11: Synthesis timings for all the benchmarks with and without Tenspiler’s tree-based optimization. Benchmark name legend in  Appendix B.
In our ablation study, we evaluate using our benchmark suites the effectiveness of the optimizations (described in Sec. 5) in making synthesis scale.
Bounded Synthesis. In this experiment, we keep the type-based filtering and tree approach while removing the incremental bounded synthesis optimizations. We start with a static tensor bound of 4 instead of the incremental approach. With this, 6 of the 12 blend benchmarks time out. In addition, benchmarks involving 2D tensors that do not time out see an average of 36.75$\times$ slowdown.
Tree Approach. For this experiment, we include type-based filtering and remove the expression tree approach for grammar filtering. We assume a fixed depth for the grammar, i.e., including all operators up to the specified depth, and increase it upon synthesis failure (starting at depth 1). Without static analysis, no assumptions are made about the operators, slice indices, variables, or constants, necessitating their synthesis. Unlike the tree approach with a fixed number of placeholders, this approach exhibits scalability issues as the number of grammar choices increases exponentially with depth. Therefore, only benchmarks with depth 1 and 2 expressions could be successfully synthesized.
As illustrated in Fig. 11, without the tree based optimization, 42 out of the total 69 benchmarks timed out. In particular, all benchmarks of the blend, blas, dspstone, makespeare, and utdsp suites timed out. For the benchmarks that succeed, the synthesis phase slows down by an average 101.55$\times$ compared to the tree-based grammar filtering approach due to the need to synthesize additional expressions in both ps and inv.
6.5 Comparison with LLMs
LLMs have shown promising results in various programming-related tasks, such as code generation, translation, and testing. However, these models suffer from a lack of formal verification of the translated code and face challenges in adapting to new DSLs or backends that are not well-represented in their training corpus.
To test the capabilities of LLMs in generating code for new or low-resource DSLs, we prompt a state-of-the-art proprietary state of the art LLM Claude Opus [1] (our evaluation using other LLMs such as GPT4 shows similar results).
We selected three backends for this experiment: MLX, a completely new DSL, Gaudi, and Gemmini, which are not well represented in the training corpus of these models.
We prompted Claude Opus to generate code for these DSLs for the linear_dodge benchmark from the blend suite. The prompt instructions are shown in  Fig. 12. In Fig. 13 we show the code generated by the LLM for the three prompts.
Upon analysis, we found that all three generated programs were incorrect. The Gemmini-generated code in  Figure 13(a) partially uses the correct APIs (mvin, mvouts), but the computation with config_ex is incorrect. For the Gaudi-generated code in Figure 13(b), the model hallucinates the TPC library, which does not exist in the actual Gaudi programming model. The MLX-generated code in Figure 13(c) has the correct call to the library function add, but the imports are incorrect, making the code non-functional. In addition to the generated code being incorrect, it is challenging to verify these outputs formally as syntactically translating the generated code to SMT-LIB is not trivial. The experiment highlights two significant challenges in generating verified code using LLMs mentioned earlier. In contrast, Tenspiler, which uses a verified lifting-based approach, can easily handle these challenges.
Figure 12: Prompts for LLM
(a) LLM generated Gemmini Code
(b) LLM generated Gaudi Code
(c) LLM generated MLX Code
Figure 13: LLM generated code for the prompt
7 Related Work
Verified Lifting. Verified lifting uses program synthesis to translate code instead of designing traditional pattern-matching compilers, and has been used across application domains [15, 5, 6, 20]. Adapting these prior compilers for translation to tensor operations is nontrivial. Tenspiler introduces a novel tensor algebra-based to make synthesis efficient, and supports a diverse set of backends.
Code Translators. Tenspiler differs from other code translation approaches. While symbolic methods like pattern-matching compilers [29] face challenges with the error-prone nature of their rules, Tenspiler uses a search-based approach to avoid these complexities. Neural techniques [31, 27], treat translation as a machine translation task but struggle to ensure correctness. In contrast, Tenspiler uses a theorem prover to guarantee semantic equivalence between the translated and source code. More recently, despite the success of LLMs in programming tasks, they are unable to translate code to unfamiliar frameworks or custom hardware ISAs. Tenspiler’s approach of searching in an Tensir and using simple rules for translation makes it easy to support new backends.
Intermediate Representations. LLVM [23], MLIR [24] and TACO’s IR [21] are examples of IRs that generate code to multiple backends. LLVM in addition can generate optimized code for various hardware targets. MLIR introduces “dialects,” allowing specific optimizations for different domains or hardware targets. Despite their versatility, LLVM and MLIR were originally designed for traditional pattern-matching compilers, posing challenges for search-based compilers due to their extensive set of operators. In contrast, Tensir is designed for expressing tensor operations to be used in search-based compilers. As discussed, Tensir enables efficient lifting, verification, and code generation.
8 Conclusions
We presented our experience in building Tenspiler, a compiler that leverages verified lifting to transpile code to leverage tensor processing infrastructures. At the core of Tenspiler is Tensir which concisely captures various tensor computations. Tenspiler efficiently translates all 69 real-world benchmarks and can generate code to be executed on 6 different software and hardware backends. The generated code achieves an average speedup of 105$\times$ for kernel and 9.65$\times$ for end-to-end execution compared to the input.
Appendix
Appendix A Backend Details
Below we discuss the details of six different target DSLs that are currently supported in Tenspiler.
  • Numpy: NumPy is a Python numerical computing library that provides support for tensor operations.
  • TensorFlow: TensorFlow is a leading open source ML framework and provides a rich set of functions for creating and manipulating tensors.
  • PyTorch: Similar to TensorFlow, PyTorch is one of the leading optimized tensor libraries for DL.
  • Apple MLX: MLX was released in December 2023 as an array framework for machine learning on Apple Silicon.
  • TPC-C for Intel Gaudi processor: Intel Gaudi processors are specialized accelerators for high-performance AI, featuring a Matrix Multiplication Engine for deep-learning operations and a programmable Tensor Processor Core (TPC) cluster, a VLIW SIMD processor, that accelerates remaining operations.
    The way to execute programs on TPC cluster is by writing those programs using TPC programming language. Intel’s SynapseAI TPC SDK includes an LLVM-based TPC-C compiler to compile those programs. Specifically, the TPC programming language, TPC-C in short, is a derivative of C99 with added data types that enable easy utilization of TPC’s SIMD capabilities. It natively supports wide vector data types to assist with programming of TPCs (for example, float64, uchar256, etc.)
    A TPC program consists of two parts: TPC kernel code and host glue code. TPC kernel code is the ISA executed by the TPC processor. The host code is executed on the host machine and specifies details of how the program input/outputs can be dynamically partitioned between the numerous TPC processors in the Gaudi processor.
    Because the blend benchmarks were not using matrix multiplication operation, we decided to ask Tenspiler to generate target programs in TPC-C. In particular, TPC-C offers several intrinsics [^8] A detailed list of intrinsics can be found online at https://docs.habana.ai/en/latest/TPC/TPC_Intrinsics_Guide/index.html for typical vector operations, such as element-wise operations, reduction operations, etc. For example, v_f32_add_b can perform element-wise vector addition of two {64xfloat} vectors. Llama benchmarks, on the other hand, were mostly using matrix multiplication operation, which is not accessible via TPC-C. These operations can only be accessed via PyTorch APIs, and hence we used PyTorch as a DSL to execute Llama benchmarks on Intel Gaudi processor.
  • Gemmini (Accelerator): Gemmini [18] is an open-source, full-stack Deep Neural Networks (DNN) accelerator generator for DNN workloads. Gemmini-generated accelerators are optimized for matrix multiplications by utilizing a systolic array in Gemmini’s architecture. Given the absence of a physical Gemmini hardware accelerator, we evaluate the programs on a simulator featuring a 16x16 systolic mesh, 256 KB scratchpad for inputs and weights, a 64 KB accumulator, and an In-order RocketChip CPU.
Table  1 shows the number of DSL APIs used in the kernel code generated by Tenspiler for all the target DSLs and for all the benchmarks used in the experiments. Note that we only count the number of DSL APIs of kernel code, not the host code (or the glue code) that is needed to launch the kernels on GPUs or accelerator devices.
Benchmark C++ Numpy TensorFlow PyTorch MLX TPC-C Gemmini
color_burn 18 5 5 5 5 21 NA
color_dodge 18 4 4 4 4 20 NA
darken_blend 18 2 2 2 2 4 NA
dissolve_blend 19 8 8 8 8 7 NA
lighten_blend 18 2 2 2 2 4 NA
linear_burn 14 2 2 2 2 6 2
linear_dodge 14 1 1 1 1 4 1
multiply_blend 14 2 2 2 2 4 NA
normal_blend_8 7 4 4 4 4 17 5
normal_blend_f 8 4 4 4 4 8 3
overlay_blend 18 12 12 12 12 32 NA
screen_blend 14 4 4 4 4 12 4
matmul 13 1 1 1 1 NA 1
rmsnorm1 6 2 2 2 2 NA 3
rmsnorm2 8 7 8 8 7 NA NA
softmax1 7 2 2 2 2 NA NA
softmax2 8 3 3 3 3 NA NA
softmax3 7 2 2 2 2 NA 2
softmax4 7 2 2 2 2 NA NA
transformer1 18 13 14 14 14 NA NA
transformer2 17 10 10 10 10 NA NA
transformer3 8 7 7 7 7 NA NA
transformer4 7 3 3 3 3 NA NA
Table 1: Lines of code for C++ inputs and the number of DSL APIs in code generated by Tenspiler for blend and Llama. (For the Llama benchmarks, we do not generate TPC-C code as most use matrix multiplication, which is not supported via TPC-C. For these benchmarks, we use PyTorch DSL for Gaudi processor.)
Appendix B Benchmark Name Mapping
Here we include the number to benchmark name mapping used in all benchmark-related plots.
Blend
  1. 1.
    color_burn
  2. 2.
    color_dodge
  3. 3.
    darken_blend
  4. 4.
    dissolve_blend
  5. 5.
    lighten_blend
  6. 6.
    linear_burn
  7. 7.
    linear_dodge
  8. 8.
    multiply_blend
  9. 9.
    normal_blend
  10. 10.
    normal_blend_f
  11. 11.
    overlay_blend
  12. 12.
    screen_blend
Llama
  1. 1.
    matmul
  2. 2.
    rmsnorm_part1
  3. 3.
    rmsnorm_part2
  4. 4.
    softmax_part1
  5. 5.
    softmax_part2
  6. 6.
    softmax_part3
  7. 7.
    softmax_part4
  8. 8.
    transformer_part1
  9. 9.
    transformer_part2
  10. 10.
    transformer_part3
  11. 11.
    transformer_part4
Blas
  1. 1.
    dot
  2. 2.
    gemv
Darknet
  1. 1.
    mag_array
  2. 2.
    matrix_add_matrix
  3. 3.
    mse_array
  4. 4.
    mult_add_into_cpu
  5. 5.
    ol_l2_cpu1
  6. 6.
    ol_l2_cpu2
  7. 7.
    scale_array
  8. 8.
    scale_matrix
  9. 9.
    sum_array
  10. 10.
    translate_array
Utdsp
  1. 1.
    fir_small
  2. 2.
    lmsfir1
  3. 3.
    lmsfir2
DSP
  1. 1.
    matadd
  2. 2.
    matscal
  3. 3.
    matsub
  4. 4.
    vadd
  5. 5.
    vcopy
  6. 6.
    vmul
  7. 7.
    vneg
  8. 8.
    voffset
  9. 9.
    vrecip
  10. 10.
    vscal
  11. 11.
    vsub
  12. 12.
    wvec
Dspstone
  1. 1.
    mat1x3
  2. 2.
    n_real_updates
Makespeare
  1. 1.
    sum_of_squares
Mathfu
  1. 1.
    diveq
  2. 2.
    diveq_sca
  3. 3.
    len
  4. 4.
    len_sq
  5. 5.
    matmul_sca
  6. 6.
    muleq
  7. 7.
    muleq_sca
  8. 8.
    negate
  9. 9.
    pluseq
  10. 10.
    subeq
  11. 11.
    subeq_sca
Simpl Array
  1. 1.
    array_inc
  2. 2.
    array_sum
  3. 3.
    cube_in_place
  4. 4.
    fourth_in_place
  5. 5.
    sum_elts
Appendix C Configurations of Hardware Devices
Table 10 presents the details of the hardware devices used for our experiments.
Device Type Configuration Software Version
CPU Intel Xeon Platinum 8380 CPU 2-socket, 40 cores per socket, 2.30 GHz 512GB mem gcc 8.3.0 -O3 Python 3.9.18 NumPy 1.26.2
GPU NVIDIA V100-PCIE-16GB PyTorch-2.1.2, TensorFlow-2.15, CUDA 12.2
GPU Apple M1 Pro with a 10-core CPU, 16-core GPU, and 16-core Neural Engine MLX 0.7
Accelerator Intel Gaudi2 processor (HL-225) with 24 Tensor Processor Cores (TPCs), 2 Matrix Multiplication Engines (MMEs) 48MB of on-chip SRAM, and 96 GB of HBM2E memories hl-gaudi2-1.11.0 -fw-45.1.1-sec-5, Ubuntu 20.02 docker, Habana PyTorch 2.1.1
Accelerator 16x16 systolic mesh, 256 KB scratchpad for inputs and weights, a 64 KB accumulator, and an In-order RocketChip CPU. riscv64 -linux-gnu-gcc 12.2.0 Chronologic VCS S-2021.09-SP1-1_Full64
Table 2: Configurations of hardware used for experiments101010 Notices & Disclaimers Performance varies by use, configuration and other factors. Learn more on the Performance Index site. Performance results are based on testing as of dates shown in configurations and may not reflect all publicly available updates. See backup for configuration details. No product or component can be absolutely secure. Intel technologies may require enabled hardware, software or service activation. © Intel Corporation. Intel, the Intel logo, and other Intel marks are trademarks of Intel Corporation or its subsidiaries. Other names and brands may be claimed as the property of others.